If AR bisects CAB , and mCAR = 4x −8 , and mCAB = 72 , find the value of x.

Answer by ikleyn(52943)

.

Since AR bisects the angle CAB,  m(CAB) = 2*m(CAR).


In other words,


    72 = 2*(4x-8).


Simplify and find x


    72 = 8x - 16

    72 + 16 = 8x

    88 = 8x

     x = 88/8 = 11.


ANSWER.  x is 11 degrees.